OpenStudy (anonymous):

Use the arc length formula and the given information to find θ. s = 7 m, r = 12 m; θ = ?

OpenStudy (jhannybean):



OpenStudy (jhannybean):

So.. Arc Length formula \[\large s = \frac{n}{360}\cdot 2 \pi r\] Givens \[\large {s =7 m \\ r = 12 m}\]\[\large n= \theta \] plug them in first. \[\large 7 =\frac{n}{360}\cdot 2 \pi (12)\] Simplify RHS \[\large 7 = \frac{n}{360}\cdot 24 \pi\] we're solving for n.an you evaluate this now? :)

OpenStudy (anonymous):

Yes thank you sooo much!

OpenStudy (jhannybean):

Noo problem, good luck!